OpenStudy (anonymous):

find the exact value of t in 5*3^t=175

OpenStudy (anonymous):

3^t=175/5 3^t=35

OpenStudy (anonymous):

3^t = 35 take the log base 3 on both sides \[t = \frac{ \log 35 }{ \log3 }\] which is \[t = \frac{ \log5+\log7 }{3 }\]

OpenStudy (anonymous):

apply log .... ss same @robz8